1.00 × 1020electrons flow through a cross section of a 4.50-mm-diameter iron wire in 6.00 s .

What is the electron drift speed?
Answer #1

q = 1.602 * 10 ^ -19 C
I = q * 10 ^ 20 / 6 = dQ/dt
I = 2.67 A
GIVEN: electron density for iron:
n = 8.5 * 10 ^ 28 q/ m^3
r = 0.0045 / 2 m
A = pi * r ^ 2

= 1.5904*10^-5
Drift velocity:
vd = I / (n * q * A)


= 1.2328*10^-05 m/s